OpenStudy (anonymous):

In some other universe, K. Pack believes he has discovered a relationship between a planets distance D from its sun and the time T it takes to orbit its sun. This formula is T=radical D^9, whre T is in Earth years and D=1 cooresponds to the distance between Earth and our sun, or 93,000,000 miles. (a) planet Eta is 35 times farther from its sun than Earth is from our sun. Estimate the number of Earth years required for Eta to orbit its sun according to K.Pack's formula. (b) write this formula with rational exponents. Help please...sooo..lost

OpenStudy (dumbcow):

A) D=35 since its 35X farther away than earth is to its sun T = sqrt(D^9) = sqrt(35^9)

OpenStudy (dumbcow):

rational exponents means as a fraction i believe sqrt is same as 1/2 power so T = 35^(9/2)

OpenStudy (anonymous):

frm where do u find such questions